# Assessing Language Comprehension in Large Language Models Using Construction Grammar

### v1

@@ Abstract
Large Language Models, despite their significant capabilities, are known to fail in surprising and unpredictable ways. Evaluating their true ‘understanding’ of language is particularly challenging due to the extensive web-scale data they are trained on. Therefore, we construct an evaluation to systematically assess natural language understanding (NLU) in LLMs by leveraging Construction Grammar (CxG), which provides insights into the meaning captured by linguistic elements known as constructions (Cxns). CxG is well-suited for this purpose because provides a theoretical basis to construct targeted evaluation sets. These datasets are carefully constructed to include examples which are unlikely to appear in pre-training data, yet intuitive and easy for humans to understand, enabling a more targeted and reliable assessment. Our experiments focus on downstream natural language inference and reasoning tasks by comparing LLMs’ understanding of the underlying meanings communicated through 8 unique Cxns with that of humans. The results show that while LLMs demonstrate some knowledge of constructional information, even the latest models including GPT-o1 struggle with abstract meanings conveyed by these Cxns, as demonstrated in cases where test sentences are dissimilar to their pre-training data. We argue that such cases provide a more accurate test of true language understanding, highlighting key limitations in LLMs’ semantic capabilities. We make our novel dataset and associated experimental data including prompts and model responses publicly available.
Assessing Language Comprehension in Large Language Models Using Construction Grammar
@@
Molly Shichman Army Research Lab Taylor Hudson Army Research Lab Claire Bonial Army Research Lab Harish Tayyar Madabushi University of Bath
@@ 1 Introduction
Understanding the language faculties of Large Language Models (LLMs) has become increasingly important with their introduction into the mainstream. LLMs have made surprising performance gains on a wide swath of NLP tasks, but continue to fail in seemingly unpredictable and idiosyncratic ways. Linguistic theory is a possible lens for understanding these failures and LLMs more generally. Indeed, analysis of transformer language models has shown that they implicitly encode substantial linguistic information (e.g., Rogers et al., 2021; Manning et al., 2020). However, there is considerable doubt about the compatibility of traditional generative linguistic theories with LLMs (Piantadosi, 2023).
This work tests model ‘understanding’ through the creation of specific evaluation tasks based in CxG (Goldberg, 1995; Croft, 2001). CxG encompasses a family of linguistic theories which posit that morphemes, words, idioms, and even schematic structures in language can be represented as simple form-meaning pairs, or constructions (Cxns), at different levels of granularity and abstraction. Because Cxns denote semantics associated with clear syntactic patterns, they provide a mechanism to develop test instances of true ‘understanding.’ Specifically, we create an NLI task where understanding the Cxn is necessary for the model to complete the task. For example, consider the following instance of the let-alone Cxn, which invokes a scale of a property (here ease of consistency), where the slot preceding “let alone” is lower on that scale than the slot after: “It is difficult enough for an individual to be consistent let alone a society.” To test if a model understands this Cxn, we convert this to an NLI problem by using the original sentence as the premise and introducing the hypothesis “It is easier for a society to be consistent than an individual.” The NLI task is then to determine if the hypothesis follows from the premise (entailment), contradicts it (contradiction), as in the case above, or if no conclusion can be drawn (neutral). Performing this task requires both knowledge of a Cxn’s meaning and reasoning over that meaning and the surrounding context.
The use of Cxns to evaluate LLM ‘understanding’ of language offers two significant advantages. First, it allows for a systematic evaluation of language understanding by leveraging the well-studied semantic characteristics of the Cxn itself, such as the contrastive scale set up by let-alone illustrated above. Second, it enables the systematic collection of test cases that are easily understood by native speakers due to familiarity with the underlying Cxn, but are relatively uncommon or do not capture statistical co-occurrence. This second advantage is particularly important, as modern LLMs are trained on web-scale corpora, making it challenging to evaluate their performance on cases they have not been previously exposed to.
1.1 Contributions
This work aims to explore the extent to which LLMs truly ‘understand’ language through the lens of CxG, specifically targeting understanding of a variety of 8 English phrasal Cxns. This study differs from previous research in several significant ways that are highlighted by the following contributions:
@@ 1.1 Contributions
We create multiple novel datasets using constructional information designed to evaluate the capabilities of LLMs.
@@
Rather than focusing on the metalinguistic task of identifying Cxns, as most prior works have done, we use downstream reasoning and the well established natural language inference (NLI) task to evaluate LLM ‘understanding’ of the underlying meaning communicated by a Cxn.
@@
We test reasoning in challenge scenarios where memorization and pattern matching from large pre-training datasets can provide less assistance.
@@
We test on the largest models currently available, including GPT-4o.
The structure of this paper is as follows: We begin with a brief introduction to CxG (§2.1) and other relevant literature (§2.2). We then introduce our experimental design and its grounding in CxG theory (§3). We detail each of three different construction-based experiments (§4, §5, and §6), along with error analysis (§7). Finally, we provide an overview of our results and their implications (§8), before concluding with possible avenues of future work (§9). Overall, our evaluations indicate that LLMs struggle to understand all the nuances of constructional meaning, particularly in situations where they must generalize meaning to examples that are not statistically likely.
@@ 2 Related Work
2.1 Construction Grammars
Although there are several somewhat differing schools of CxG, most emphasize the usage-based nature of language acquisition as well as the fundamental unit of language: the Cxn. Cxns are pairings of form and meaning at all levels of language; thus, there are morphological Cxns, lexical Cxns, and phrasal Cxns. In this paper, we focus on phrasal Cxns. In the CxG account, speakers use domain-general cognitive processes to incrementally generalize over frequent Cxns, beginning with single-word Cxns, to novel usages, arriving at the ability to interpret even rare and previously unseen instances of a Cxn. Therefore, CxG provides an effective method of evaluating the true generalization capabilities of LLMs. We provide a more detailed description of CxG tenets in Appendix A.
@@ 2.1 Construction Grammars
2.2 CxGs and LLMs
Starting with CxGBert (Tayyar Madabushi et al., 2020), there has been substantial past work on probing language models’ understanding of Cxns.111These works typically take a set of Cxns from CxG theoretical literature for their experiments. See Dunn (2017, 2023) for related work which attempts to discover Cxns from embedding spaces using a bottom-up approach. These works have typically focused on either a single Cxn or a handful of Cxns, like aann (Mahowald, 2023; Chronis et al., 2023), comparative-correlative (Weissweiler et al., 2022), and more schematic phrasal Cxns (Li et al., 2022; Veenboer and Bloem, 2023). Tseng et al. (2022) focus on Cxns in Taiwanese Mandarin, a notable exception to the works above which focus on English. Zhou et al. (2024) show that LLMs fail to adequately distinguish examples of the causal-excess Cxn from similar patterns, and introduce NLI as a proxy task for understanding constructions, though their results are limited to the Causal-Excess construction. Bonial and Tayyar Madabushi (2024) provide a corpus of 10 Cxns across varying levels of schematicity, showing that GPT-4 is not proficient at grouping sentences according to constructional similarity. Of particular relevance to our work is Weissweiler et al. (2022), who define a set of bias and calibration tests for the comparative-correlative, which we adapt as challenge reasoning tasks for the comparative-correlative, let-alone, and way-manner Cxns.
3 Theory-Informed Experimental Design
CxG’s usage-based account of the development of a grammar is particularly relevant to the evaluation of LLMs because language models learn through memorization, frequency and collocational properties. However, it is not clear the extent to which LLMs can recognize and generalize over the formal properties of language in order to ascribe shared semantics to novel usages (e.g., resultative: “The jackhammer pounded us deaf”). Indeed, such an ability to generalise would be a key indicator of ‘understanding.’ To explore this question, we leverage the CoGS corpus Bonial and Tayyar Madabushi (2024), which is a collection of about 500 corpus instances, roughly balanced across 10 different Cxns.
CoGs consists a limited set of carefully curated Cxns chosen for their broad coverage of the English language. These Cxns collectively represent a significant portion of English usage and provide an effective basis for evaluating LLMs in both common and uncommon linguistic scenarios. Our primary goal is not to test LLMs on specific Cxns, but rather to use constructional information to design targeted test cases that are easy for humans to understand. This allows a more precise evaluation of LLMs, resulting in the number of Cxns we use being less critical to allowing us to generalize our findings to the English language as a whole. Of the Cxns in CoGS, we construct our datasets with 8 of the 10 Cxns, shown in Table 1. In designing our dataset, we faced a trade-off between quality and quantity. While generating a large number of test samples semi-automatically was an option, we prioritized creating a smaller, carefully curated set of high-quality examples. This not only allowed for meticulous manual verification of the curated data but also allowed us to manually validate the results, which would be infeasible with a larger dataset. By minimizing biases and errors in both dataset creation and evaluation we ensure the reliability of our findings, even if it comes at the cost of a reduced number of examples.
4 Experiment 1: Constructional NLI
4.1 Datasets
We take Natural Language Inference (NLI) as a good starting point for investigating LLM capabilities regarding the semantics of Cxns, following (Zhou et al., 2024). NLI tasks present a premise, taken to be true, and a hypothesis, which can be entailed by the premise, contradicted by the premise, or neutral (neither contradicted nor entailed). Thus, by targeting entailment and contradiction, NLI effectively probes for a functional understanding of a text. NLI is also a widely accepted task for testing NLU in the NLP community, as evidenced by the many shared evaluations ranging from the general evaluation of distributional semantic models Marelli et al. (2014) to the evaluation of LLMs for biomedical applications Lee et al. (2024).
In particular, we introduce an NLI dataset which specifically probes the semantics of our Cxns. We do this by templatically generating hypotheses which make use of certain slots within a Cxn, and verifying the resultant sentences manually to ensure quality. These template-based hypotheses have predictable relations to the premises. For example, consider the premise: "He hammered the metal flat". This is a resultative Cxn, which has the meaning of an action which causes a change in state. In this case, the action verb hammered leads to the metal to have a resulting state of flat. Regarding the syntactic form of a phrasal Cxn, we can use a constructional template to define the slots that are filled by a given Cxn. A general template for the resultative is shown in Example 4.1, and its application to the above sentence in Example 4.1.
. [SBJ1 [V2 OBJ3 ADJ4 ]VP]5
. [He1 [hammered2 the metal3 flat4 ]VP]5
Given constructional examples like those in Example 4.1, our goal is to produce NLI tuples that consistently target the construction’s meaning. We do this by manipulating the slots in the Cxn templatically to produce hypotheses with consistent relations to the premise. For instance, consider the hypothesis "The hammering did not cause the metal to become flat." This is clearly a contradiction to the above premise and is directly in conflict with the meaning of the Cxn. We can produce this hypothesis, and similar hypotheses, by following the template in Example 4.1. We can see this template reflecting the above hypothesis in Example 4.1.
. [The [V]2-ing did not cause [OBJ]3 to become [ADJ]4].
. [The [hammer]2-ing did not cause [the metal]3 to become [flat]4].
For each Cxn, we create a template to generate NLI hypotheses that target constructional meaning.222There are many ways to create NLI hypotheses beyond the templates we utilize for our hypotheses. We also create a more freeform NLI dataset given our constructional premises, finding similar results. These templates produce hypotheses that have predictable relations, and are directly tied to the semantics of the Cxn. See Appendix B for details. See Appendix E for the NLI templates for each Cxn, along with examples. Overall, our process for creating the constructional NLI dataset can be summarized as:
@@ 4.1 Datasets
Extract natural constructional examples from CoGS Bonial and Tayyar Madabushi (2024).
@@
Create general templates for NLI hypotheses for each Cxn.
@@
All templatically generated hypotheses and relations were validated by an author. Once the dataset was created, a second author annotated the relations, achieving an Inter-Annotator Agreement (IAA) of 90% agreement. Descriptive statistics for this dataset along with all other datasets can be found in Appendix C.

### v2

# Beyond Memorization: Assessing Semantic Generalization in Large Language Models Using Phrasal Constructions

@@ Abstract
The web-scale of pretraining data has created an important evaluation challenge: to disentangle linguistic competence on cases well-represented in pretraining data from generalization to out-of-domain language, specifically the dynamic, real-world instances less common in pretraining data. To this end, we construct a diagnostic evaluation to systematically assess natural language understanding in LLMs by leveraging Construction Grammar (CxG). CxG provides a psycholinguistically grounded framework for testing generalization, as it explicitly links syntactic forms to abstract, non-lexical meanings. Our novel inference evaluation dataset consists of English phrasal constructions, for which speakers are known to be able to abstract over commonplace instantiations in order to understand and produce creative instantiations. Our evaluation dataset uses CxG to evaluate two central questions: first, if models can ‘understand’ the semantics of sentences for instances that are likely to appear in pretraining data less often, but are intuitive and easy for people to understand. Second, if LLMs can deploy the appropriate constructional semantics given constructions that are syntactically identical but with divergent meanings. Our results demonstrate that state-of-the-art models, including GPT-o1, exhibit a performance drop of over 40% on our second task, revealing a failure to generalize over syntactically identical forms to arrive at distinct constructional meanings in the way humans do. We make our novel dataset and associated experimental data, including prompts and model responses, publicly available.111https://anonymous.4open.science/r/beyond-memorization-B82B
Beyond Memorization: Assessing Semantic Generalization in Large Language Models Using Phrasal Constructions
@@
Mollie Shichman Army Research Lab Taylor Hudson Army Research Lab Claire Bonial Army Research Lab Harish Tayyar Madabushi University of Bath
@@ 1 Introduction
| Model |
| GPT-4o | 0.88 | 0.58 | |||||
| GPT-o1 | 0.90 | 0.46 | |||||
| Llama 3 70B | 0.74 | 0.52 | |||||
| Human | 0.90 | 0.83 | |||||
Understanding the extent to which Large Language Models (LLMs) generalize from relatively frequent phenomena well-represented in pretraining data to creative, novel usages of language has important implications for LLM development. Identifying the precise nature and limits of LLM generalization can inform decisions about architectures and training regimes Li et al. (2023); Zhang et al. (2023). This becomes especially relevant as models move toward ‘reasoning’-based systems and the inevitable widespread deployment of AI agents DeepSeek-AI et al. (2025). Identifying failure patterns will enable targeted improvements at different stages of development. However, testing LLMs’ ability to generalize is particularly challenging because they are trained on vast web-scale data Lu et al. (2024). Even if pretraining datasets were fully accessible, ensuring that a test example is truly independent would remain nontrivial, as a model may not have encountered that instance but could have been exposed to related cases that provide indirect information Tayyar Madabushi et al. (2025).
| Exp 1 Dataset | Exp 2 Dataset | |
| CxNLI Premise | I brushed my hair smooth. | A famous emperor buried scholars alive. |
| Entrenched Variant | I made my hair smooth (by brushing). | NA: no entrenched variants for Exp2 Cxns |
| Cxn Form | NP1 V NP2 ADJ | NP1 V NP2 ADJ |
| Cxn Meaning | NP1’s action of V causes NP2 to become ADJ | NP2 is ADJ during NP1’s action of V |
| CxNLI Hypothesis | My brushing caused my hair to be smooth. | Burying the scholars caused them to be alive. |
| CxNLI Relation | Entailment | Contradiction |
Therefore, this work introduces a novel evaluation dataset grounded in the theory of Construction Grammar (CxG) (Goldberg, 1995; Croft, 2001) (see Appendix A for an overview of CxG). Specifically, we focus on phrasal constructions (Cxns) because of the body of psycholinguistic research demonstrating that speakers are able to abstract over syntactic slots of these Cxns in order to interpret and produce creative and novel Cxn instantiations Johnson and Goldberg (2013); Tomasello (2003). Speakers recognize the syntactic structures of a familiar Cxn in order to interpret the meaning, despite the fact that the speaker may have never encountered that set of lexical items within the Cxn. The Exp 1 column of Table 2 provides an example premise where our human annotations show that people can easily recognize the (Resultative) constructional semantics despite the fact that the verb (“brush”) is relatively atypical in this Cxn, which is realized with a relatively limited set of verbs (frequently “make”). Additionally, speakers can balance knowledge of lexical semantics against the constructional semantics in order to distinguish Cxns that are syntactically identical but have different meanings. The Exp 2 column of Table 2 provides an example premise involving a (Depictive) Cxn that is syntactically identical to the Resultative, but has a divergent meaning, as evidenced by the fact that a templatically similar hypothesis holds the opposite relation.
In addition to providing experimentally validated explanations of human language acquisition and use, CxG is uniquely suited to evaluating whether LLMs primarily derive meaning from the composition of lexical meanings (as would be the view of Generative Grammar Chomsky (2014a)), or if the recognition of particular syntactic structures can cue constructional meaning (this would be evidenced by strong performance on Exp1), or finally, if LLMs can balance both lexical meaning and constructional meaning together to recognize constructional meaning while distinguishing between syntactically identical constructions based on lexical semantics (this would be evidenced by strong performance on Exp 2).
We evaluate whether or not models can generalize knowledge of highly frequent Cxns of English to creative instantiations of those Cxns with lexical items unlikely to have been encountered within that structure in pretraining data (Exp 1), and we evaluate model ability to recognize when the lexical items are so distinct that this cues a different Cxn with a different meaning (Exp 2).
To investigate LLM generalization, we conduct two experiments (described in §3) leveraging the Natural Language Inference (NLI) task. We show a summary of results in Table 1: all models lag behind human performance in multiple experimental settings. Exp 1 (§4) shows us that some aspects of constructional semantics are ascribed adequately for success on an NLI task; however, Exp 2 (§5) shows that even state-of-the-art reasoning models like GPT-4o and GPT-o1 show a significant performance drop when the models are asked to ascribe distinct constructional semantics to syntactically identical Cxns. In combination, our results and error analysis (§6) highlight key differences between human and model linguistic capabilities (§7 and §8). This study differs from previous research in significant ways highlighted by the following contributions:
@@
We create a manually-validated, diagnostic evaluation dataset of 534 NLI triples testing if LLMs ascribe the appropriate semantics to phrasal Cxns.
@@
Rather than focusing on the metalinguistic task of identifying Cxns, as most prior works have done, we use the well-established NLI task to evaluate LLM ‘understanding’ of the underlying meaning communicated by a Cxn.
@@
We test on a variety of common English phrasal Cxns instantiated by relatively unexpected words, thereby testing the ability of models to perform understanding tasks without the aid of memorization and pattern matching from large pretraining datasets.
@@
We test some of the largest models currently available, including GPT-o1 and Llama 3 70B.
@@ 2 Related Work
The evaluation of LLMs with datasets that consider pretraining data has largely focused on identifying and mitigating test set data leaks Balloccu et al. (2024); Sainz et al. (2023a). Examples of this include work by Golchin and Surdeanu (2024) and Sainz et al. (2023b) who identify data contamination using prompting. Furthermore, considerable effort has been directed toward designing datasets that minimize such leakage Zhou et al. (2025). While this is an important concern, preventing direct data leakage does not eliminate the possibility that models can infer answers using related information from pretraining data Tayyar Madabushi et al. (2025).
Another relevant line of research involves counterfactual reasoning, where standard rules of the world are slightly altered. By design, counterfactuals provide an effective way to test LLMs in scenarios where pretraining data offers little advantage, and it has been shown that the performance of LLMs does in fact drop in these cases Wu et al. (2024); Lewis and Mitchell (2024).
Specific to CxGs, starting with CxGBert (Tayyar Madabushi et al., 2020), there has been substantial past work on probing language models’ understanding of Cxns. These works have typically focused on either a single Cxn or a handful of Cxns, like aann (Mahowald, 2023; Chronis et al., 2023), comparative-correlative (Weissweiler et al., 2022), and more schematic phrasal Cxns (Li et al., 2022; Veenboer and Bloem, 2023). Tseng et al. (2022) focus on Cxns in Taiwanese Mandarin, a notable exception to the works above which focus on English. Zhou et al. (2024) introduce NLI as a proxy task for understanding Cxns, though their results are limited to the Causal-Excess and related Cxns. We expand on the use of NLI to a broad set of new Cxns, and create NLI examples which utilize both entrenched (Exp 1) and syntactically-identical (Exp 2) Cxns.
To our knowledge, no prior work has explored the use of cognitive linguistic principles to generate human-generalizable datasets for assessing the generalization capabilities of LLMs.
3 Experimental Design
We conduct experiments exploring two questions:
Research Question (RQ) 1: To what extent can models generalize constructional semantics to relatively infrequent instantiations of common Cxns? Exp 1 uses NLI to evaluate understanding of constructional semantics in cases where high-frequency constructional templates are instantiated by words not commonly found within that Cxn. We select 8 Cxns that are roughly balanced across two types: argument structure Cxns with no fixed words but clear syntactic slots (e.g., caused-motion in Table 3) and phrasal Cxns with two or more fixed words that clearly identify the Cxn (e.g., let-alone in Table 3). To evaluate if the appropriate constructional semantics are associated with these Cxns, we create a novel NLI dataset, where premises are derived from corpus instances of Cxns and an understanding of constructional semantics is required to determine entailment.
RQ 2: To what extent can models distinguish the semantics of Cxns that are syntactically identical, but have different meanings? Exp 2 uses NLI to evaluate abstraction of distinct constructional semantics given identical syntactic phrasings. We select five Cxns that are syntactically identical to the five argument structure Cxns of Exp 1. We create a second set of NLI instances again using corpus instances of the five semantically distinct, but syntactically identical Cxns.
Parallel to psycholinguistic research on analogical extension of Cxns Bybee (2010), we hypothesize that the frequency and entrenchment of the Cxn contribute to model ability to understand the constructional semantics. There is abundant corpus linguistic data from both web and even child language indicating that the 5 argument structure Cxns tested in Exp 1, Causative-with, Caused-Motion, Conative, Intransitive Motion, Resultative, are some of the earliest acquired and most frequently used Cxns in the English language Hoffmann (2022); Tomasello (2003); Gries and Stefanowitsch (2004). Although we lack CxG-annotated resources and access to pretraining data to calculate the precise frequencies of the Cxns tested in Exp 2 (see Table 5), we can safely assume that these Cxns are less frequent in the language.222We do not argue that the type frequencies of all Exp 2 Cxns are low (e.g., Intransitive has high type frequency), rather that the specific sub-Cxns of Exp 2 (e.g., Intransitive+”at”-Locative) are lower than type frequencies of argument structure Cxns of Exp 1. Most importantly, while the Cxns of Exp 1 are generally instantiated by a more limited set of verbs, the Cxns of Exp 2, such as the Depictive and Locative, can co-occur felicitously with any verb. As a result, there is no single, entrenched variant of the Cxns in Exp 2 Gries and Stefanowitsch (2004). Thus, the Cxns of Exp 1 have higher type frequency (the frequency of, for example, the Resultative overall) and have at least one entrenched variant with high token frequency (the frequency of the Resultative with “make”). Higher entrenchment of one variant means that there is a strong lexico-syntactic signature associated with a specific meaning of a Cxn. Lower entrenchment indicates that there is more variation in the lexico-syntactic features and more variation in how identical lexico-syntactic features are associated with several meanings. Thus, greater entrenchment may provide critical priors for the model to generalize constructional semantics to novel instantiations, whereas these priors may not be available for the Exp 2 Cxns, which lack any entrenched, high token-frequency exemplar.
@@ 3 Experimental Design
4 Experiment 1: NLI for Constructional Semantics
4.1 Dataset
In Exp 1, we leverage the CoGS corpus Bonial and Tayyar Madabushi (2024b), which is a collection of about 500 corpus instances, roughly balanced across 10 different Cxns. CoGs consists of carefully curated Cxns chosen for their broad coverage of the basic phrasal Cxns of English Hoffmann (2022). The Cxn types collectively represent a significant portion of English usage and provide an effective basis for evaluating LLMs on high-frequency Cxns (such as the Caused-Motion), where instantiated with creative words.
Broadly, CoGS consists of Cxns of two types: argument structure Cxns, which involve no fixed word forms but have been shown to be the most common Cxns of English as well as other languages Goldberg (1995); and phrasal Cxns with multiple fixed words. The latter Cxns are more easily recognizable to LLMs given fixed words cueing that Cxn Bonial and Tayyar Madabushi (2024a). We construct our datasets with 8 of the 10 Cxns in CoGS, shown in Table 3.
Overall, our process for creating the constructional NLI dataset can be summarized as:
@@ 4.1 Dataset
Extract corpus Cxn examples from CoGS.
@@
Create general templates for NLI hypotheses for each Cxn type.
@@
We explain this process through the following example, beginning with the premise: “He hammered the metal flat.” This is a resultative Cxn, which has the meaning of an action causing a change in state. In this case, the action verb hammered leads to the metal to have a resulting state of flat. Regarding the syntactic form of a phrasal Cxn, we can use a constructional template to define the syntactic nature of the slots that are filled by a given Cxn. A general template for the resultative is shown in Example 4.1, and its application to the above sentence in Example 4.1.
. [SBJ1 [V2 OBJ3 ADJ4 ]VP]5
. [He1 [hammered2 the metal3 flat4 ]VP]5
Given constructional examples like those in Example 4.1, our goal is to produce NLI tuples that consistently target the Cxn’s meaning. We do this by manipulating the slots in the Cxn templatically to produce hypotheses with consistent relations to the premise. For instance, consider the hypothesis “The hammering did not cause the metal to become flat.” This is clearly a contradiction to the above premise and is directly in conflict with the meaning of the Cxn. We can produce this hypothesis, and similar contradicted hypotheses, by following the template in Example 4.1, instantiated with our current example in Example 4.1.
. [The [V]2-ing did not cause [OBJ]3 to become [ADJ]4].
. [The [hammer]2-ing did not cause [the metal]3 to become [flat]4].
Thus, for each Cxn, we create a template to generate NLI hypotheses that target constructional meaning. See Appendix D for the NLI templates for each Cxn (Table 8), along with examples (Table 9).
Manual Verification We manually validate every test instance of our dataset with double or triple annotation and measure human agreement (ranging from 78-90%) to ensure the robustness of our claim that people are able interpret the specific Cxn instances that we present to the LLMs. Once the dataset was created, a second and sometimes third author annotated the relations and hypotheses were amended until achieving an Inter-Annotator Agreement (IAA) of 90%. Thus, if we take the original author’s assigned relation to be the gold standard, then native speaker accuracy on the NLI task is 90%. The final Exp 1 “CxNLI” dataset totals 435 triples. Descriptive statistics for this dataset along with all other datasets can be found in Appendix B (Table 7).
4.2 Formalism and Task Design
We define a Cxn, , to be a pairing of a constructional schema (form), , and a semantic interpretation (meaning), . For example, for the Resultative Cxn has the schema () NP V NP ADJ and Meaning () ‘The action of the Verb causes the Object to enter the state described by the Adjective.’
Our method of evaluating models’ ability to ‘understand’ this Cxn begins with a premise sentence, , that is an instance of a given Cxn (). We then generate a hypothesis, , by applying a pre-defined hypothesis template, . This template is a function that takes the premise as input, extracts its relevant components, and generates a new sentence . The template is designed to probe the Cxn’s core meaning, , in a way that produces a predictable NLI label, . Therefore, the hypothesis is generated as: . For example, to generate a contradiction () for a Resultative premise:
Premise (): “He hammered the metal flat.”
Hypothesis (): “The hammering did not cause the metal to become flat.”
Resulting Tuple: .
The goal of this experiment is to assess if models can correctly classify these NLI tuples. High accuracy on this task indicates that a model has learned the fundamental association between a syntactic schema and its meaning , even when instantiated with creative or atypical words.